OpenStudy (anonymous):

HELP PLEASE!!!! I need help understanding how to do piecewise functions: A human chain is formed by 60 people standing with their arms outstretched, each holding the hand of the person on either side. The first 30 people in the chain have arm spans of 6 feet. The next 30 people have arm spans of 5.5 feet. At the word "go!" the first person squeezes the hand of the second person. Then the second person squeezes the hand of the third, and so on. Assume that each person takes 1/3 second to pass along the signal. Which piecewise function correctly models the distance that the signal travels in t seconds?

OpenStudy (anonymous):

The problem is simplified into multiple choice answers as well: A) {6t if 0<=t<=10, 5.5(t-10)+60 if 10<t<=20 B){18t if 0<=t<=30, 16.5(t-30)+180 if 30<t<=60 C){18t if 0<t<15, 16.5(t-5)+180 if 15<=t<=20 D){18t if 0<=t<=10, 16.5(t-10)+180 if 10<t<=20

OpenStudy (lurker):

hey okay

OpenStudy (lurker):

so we got 6 feet of signal travelling every 1/3rd second 30 times then 5.5 feet of signal travel every 1/3rd second 30 times

OpenStudy (lurker):

in other words 6*3=18 feet a second for a distance of 6*30 and then 5.5*3 = 16.5 feet a second for a distance of 5.5*30

OpenStudy (lurker):